# Extending the SAREF4ENER Ontology with Flexibility Based on FlexOffers

### v2

@@ III-C Constraints Encoding
Here, we show how we encode the FO constraints into the dCO ontology: for those constraints, we will use the same notation we used in Section II. The only properties in SAREF4ENER that can be used for this are s4ener:EnergyMin and s4ener:EnergyMax: it is possible to represent SFOs and TECFOs by mapping to them $emin_{t}$ and $emax_{t}$ for slice constraints, and $TE_{min}$ and $TE_{max}$ for TECs. Further information on the FO mapping to SAREF4ENER can be found in Tables III and IV in Appendix -B.
Regarding DFOs, we show how we encoded the constraints. Consider the running example, and the DFO represented in Figure 2(c). The slices are represented by polygons, which are intersections of hyperplanes: thus, we can write these polygons as sets of inequalities over $x$ and $y$. We can then build a matrix, with as many rows as the number of inequalities defining the polygon: in this case, the number of rows would be six. As the inequalities are linear, they can be written as $ax+by\leq c$: the corresponding row of the matrix will then be $[a\;b\;c]$. This is called H-representation of the DFO [24]. We show below the constraints and matrix representation of the fourth slice. The property DependencyEnergyConstraintList encodes the matrix: details are provided in Appendix -E.
$y$ $\geq 0.309$
$y$ $\leq 0.442$
$x$ $\geq 0.972$
$x$ $\leq 1.246$
$0.088\cdot x+y$ $\geq 0.41$
$0.088\cdot x+y$ $\leq 0.537$
@@
Finally, UFOs are described by the functions $\{f_{1},\ldots,f_{T}\}$ mentioned in Section II-F. Describing a generic function is nontrivial; however, for the devices supported by FOs, we noticed that they can be written as $f_{t}=\min\{p_{1},\ldots,p_{N_{t}}\}$, where each $p_{i}$ is a polynomial defined in the range of possible energy values. In the running example, a representation of $f_{1}$ and $f_{2}$ can be seen in Figure 4: those functions are, however, represented on the $y$ axis rather than the $x$ axis, because FOs represent energy over the $y$ axis. We can see that for $t=1$ the function is the constant polynomial $f_{1}=1$, while for $t=2$ we have $f_{2}=\min\{p_{1},p_{2},p_{3}\}$, where $p_{1}=1$, $p_{2}=66.67\cdot x-20.6$, and $p_{3}=-66.67\cdot x+29.467$. Once the polynomials have been determined, they are encoded as a list of lists of coefficients, with the terms corresponding to each degree in increasing order. For example, since $f_{1}=1$, the UFO constraint for $t=1$ is encoded as $[1]$. On the other hand, for $t=2$, the list will be $[[1],[-20.6,66.67],[29.467,-66.67]]$. Details are provided in Appendix -F.
0011110.340.340.340.340.360.360.360.360.380.380.380.38Probabilityt=1Energy,t=1000.50.50.50.511110.350.350.350.350.40.40.40.4Probabilityt=2E,t=2
@@ IV Evaluation
In Section II-G, we mentioned that SAREF4ENER is an ontology that allows for a simple description of energy flexibility. FOs, on the other hand, is a model that represents flexibility in mathematical terms; despite having two different purposes, a parallel can be drawn between SAREF4ENER and FO representations, in that they refer to energy profiles and time slices. However, this parallel highlights a crucial weakness of SAREF4ENER in describing energy flexibility. Specifically, SAREF4ENER is unable to model many necessary attributes for FOs: in the main FO message, only Stateandthe attributes defining slice and total energy constraints can be mapped to SAREF4ENER. It is therefore possible to map SFO and TECFO constraints into SAREF4ENER. However, as mentioned in Section III-A, dependency constraints and UFOs cannot be mapped to SAREF4ENER. On the other hand, the dCO flexibility module has been built in order to provide full support for all types of FOs. This means not only that dCO can represent flexibility through more advanced constraints, but also that it allows for communication via a FO message. This way, by using dCO, developers have the possibility to design applications able to exploit the whole range of benefits coming from FO modeling, i.e., advanced and more accurate flexibility representation, support for more advanced devices, and really fast optimization and (dis)aggregation of flexibility. For example, FOs allow to aggregate, optimize and disaggregate up to $2\cdot 10^{6}$ of loads for 96 time units time horizons [15]. In comparison, an exact approach described in [15] can only optimize up to 500 devices for that time horizon, while other known exact approaches cited in [15] fail already to perform optimization for more than 6 time units.
@@
The domOS project demonstrated dCO feasibility in real-world case studies in Denmark, France, and Switzerland: all functionalities, e.g., querying building descriptions, are operative, and interoperability between smart buildings and services is achieved. The energy services using dCO cut down energy consumption and smoothened load consumption curves for 124 individual customers and 50 days. The cumulative amount of flexibility (load shedding) was several mWh.
@@
Regarding flexibility representation, SAREF4ENER is able at most to accurately represent flexibility for devices that only need this type of constraint. This is the case with wet devices, such as dishwashers and washing machines, as shown in [11]. These devices have a load with a fixed power/amount profile, which is only shifted in time. However, for other devices, SFOs/TECFOs are outclassed by DFOs in terms of accuracy in flexibility representation, as we can see in previous literature. The ability to retain flexibility for a model is calculated by the economic metric [19], defined as the amount of profit that can be obtained by optimizing flexibility with the objective of cost reduction. Specifically, for the charge and discharge process of batteries, DFOs retain $61$% of the flexibility, against the $38$% from TECFOs and $10$% from SFOs [19]. UFOs can perform even better, retaining up to $66.4$% of flexibility, which goes up to $86.8$% of flexibility when imbalance penalties for violating constraints are high (high penalty case) [21]. Regarding EVs, flexibility modeling via DFOs is more accurate than SFOs and TECFOs, as slice constraints and TECs are only a subset of DFO constraints, and therefore the amount of flexibility retained by DFOs is higher: yet, in [21], we see that UFOs can retain $92$% of flexibility, against the $77.3$% retained by DFOs. For the high penalty case, UFOs retain $87.5$% of flexibility, against the $61.8$% from DFOs, which still outperform SFOs and TECFOs. Finally, regarding heat pumps, the comparison between DFOs and SFOs has been done in [24], showing that DFOs can retain more than three times the amount of flexibility from SFOs: in [15] DFOs for heat pumps have been further improved, showing that they can retain up to $98.9$% of flexibility according to the economic metric. The ability for dCO to model the significant amounts of flexibility for these devices ensures that developers can create applications that allow for an accurate description of flexibility for batteries, EVs and heat pumps. Table I summarizes this, showing which use cases are supported by the dCO flexibility module and SAREF4ENER, respectively.
@@
TABLE I: Use cases supported.
Furthermore, FOs are extended over time to represent more use cases and flexibility properties: an example is their capability of modeling different energy vectors from electricity, such as heat [15]. This has been our rationale to extend SAREF4ENER to represent both FOs in general and, more specifically, DFO and UFO constraints: the resulting ontology can represent all the advanced use cases presented in this paper, as well as provide a solid foundation for future ontology extensions based on new types of FOs.
In order to validate dCO, we also used the OntOlogy Pitfall Scanner! (OOPS!) tool [^4] https://oops.linkeddata.es/ to evaluate it. OOPS! [25] is a tool, well-known in the Semantic Web community, designed with the purpose of identifying pitfalls and design errors in ontologies. There are three levels of importance in pitfalls identified by OOPS!. Critical pitfalls need to be resolved, as they may affect the ontology consistency, reasoning, and applicability. Important pitfalls would improve the ontology’s robustness if fixed, but they do not undermine the ontology’s function. Minor pitfalls would improve the ontology’s readability if fixed, but do not pose problems for the ontology. OOPS! identified quite a few minor and important pitfalls, but only 11 critical ones. However, the critical pitfalls were all of the type Defining multiple domains or ranges in properties. In our case, it was not only valid, but also necessary for properties to have multiple domains and ranges to capture the complexity of our flexibility scenario. A simple example of this is the dco:hasEnergyConstraintList property, which can have up to one s4ener:EnergyMin attribute and up to one s4ener:EnergyMax attribute. For this reason, we intentionally designed our ontology with this feature, and therefore none of the critical pitfalls identified by the OOPS! tool undermines the function of the ontology, rather, they are exactly as they should be. Thus, the ontology meets the checked requirements.
Regarding the integration of this ontology with existing standards or protocols, our ontology allows for better flexibility modeling, being able to capture more complex energy constraints and integrate FOs. Our proposed ontology is also compatible with the Common Information Model (CIM) [^5] https://www.entsoe.eu/digital/common-information-model/, which is the standard from the International Electrotechnical Commission for electric power transmission and distribution, and is capable of representing more complex energy constraints than CIM. Finally, FOs integrate with EEBus [^6] https://www.eebus.org/, another energy data standard that builds on SAREF4ENER and only has simple flexibility models.
@@ V Conclusions and Future Work
Acknowledgements
This work was supported by the domOS and FEVER projects, funded under the Horizon 2020 programme, GAs 864537 and 894240 respectively, Flexible Energy Denmark, funded by Innovation Fund Denmark (Case No. 8090-00069B), and CLIMETransition, funded under the EU Next Generation programme (CUP F53D23009450001).

### v3

@@ III-C Constraints Encoding
Here, we show how we encode the FO constraints into the dCO ontology: for those constraints, we will use the same notation we used in Section II. As we will see in Section III-D, the only properties in SAREF4ENER that can be used for this are s4ener:EnergyMin and s4ener:EnergyMax: it is possible to represent SFOs and TECFOs by mapping to them $emin_{t}$ and $emax_{t}$ for slice constraints, and $TE_{min}$ and $TE_{max}$ for TECs.
Regarding DFOs, we show how we encoded the constraints. Consider the running example, and the DFO represented in Figure 2(c). The slices are represented by polygons, which are intersections of hyperplanes: thus, we can write these polygons as sets of inequalities over $x$ and $y$. We can then build a matrix, with as many rows as the number of inequalities defining the polygon: in this case, the number of rows would be six. As the inequalities are linear, they can be written as $ax+by\leq c$: the corresponding row of the matrix will then be [abc][a\;b\;c]. This is called H-representation of the DFO [DBLP:conf/eenergy/SiksnysP16]. We show below the constraints and matrix representation of the fourth slice. The property DependencyEnergyConstraintList encodes the matrix.
[…]
$088\cdot x+y 0$$\geq 41$
$088\cdot x+y 0$$\leq 537$
@@
Finally, UFOs are described by the functions $\{f_{1},\ldots,f_{T}\}$ mentioned in Section II-F. Describing a generic function is nontrivial; however, for the devices supported by FOs, we noticed that they can be written as $f_{t}=\min\{p_{1},\ldots,p_{N_{t}}\}$, where each $p_{i}$ is a polynomial defined in the range of possible energy values. In the running example, a representation of $f_{1}$ and $f_{2}$ can be seen in Figure 5: those functions are, however, represented on the $y$ axis rather than the $x$ axis, because FOs represent energy over the $y$ axis. We can see that for $t=1$ the function is the constant polynomial $f_{1}=1$, while for $t=2$ we have $f_{2}=\min\{p_{1},p_{2},p_{3}\}$, where $p_{1}=1$, $p_{2}=66.67\cdot x-20.6$, and $p_{3}=-66.67\cdot x+29.467$. Once the polynomials have been determined, they are encoded as a list of lists of coefficients, with the terms corresponding to each degree in increasing order. For example, since $f_{1}=1$, the UFO constraint for $t=1$ is encoded as $[1]$. On the other hand, for $t=2$, the list will be [[1],[20.6,66.67],[29.467,66.67]][[1],[-20.6,66.67],[29.467,-66.67]].
0110.340.340.360.360.380.38Probabilityt=1Energy,t=100.50.5110.350.350.40.4Probabilityt=2E,t=2
@@
III-D FO and SAREF4ENER attribute comparison
In this section we map all the FO message attributes into SAREF4ENER, whenever possible, to observe which FO attributes are present in SAREF4ENER and which ones are missing. Table I and Table II show this mapping, describing, for each attribute of the FO message, whether it is mandatory or not, whether it exists in SAREF or SAREF4ENER or not, what is the corresponding class in SAREF/SAREF4ENER, and notes for additional information. More precisely, Table II shows the mapping for the main FO message attributes and slice and time constraints, while Table II shows the mapping for the more general constraints, such as TEC, dependency, and uncertainty. We can see that, among the core FO message attributes, only the attribute State has a direct counterpart in SAREF4ENER, PowerSequenceState, while other attributes are not directly supported. Duration and slice constraints can be mapped to SAREF4ENER, while SAREF has a partial match for PriceConstraint, although it does not allow specification of minimum and maximum prices. FO schedules can also be mapped to SAREF4ENER, except for price attributes, and the same applies to TECs. However, dependency and uncertainty constraints cannot be mapped to SAREF4ENER, as shown in the lower part of the table. Consequently, SAREF4ENER is not capable of representing DFOs and UFOs, and thus cannot benefit from the accurate flexibility representation given by these advanced types of FOs.
@@ IV Evaluation
This section shows the evaluation of this ontology. Specifically, how it has been demonstrated in real life and integrates with the current standards, how it has been validated, and how it compares with the current state of the art.
IV-A Real-life demonstration and standard integration
@@
The EU project in which dCO was developed demonstrated its feasibility through real-world case studies conducted in Denmark, France, and Switzerland. All functionalities, e.g., querying building descriptions, are operational, and interoperability between smart buildings and services has been successfully achieved. dCO supports the flexibility services developed in this project.
The Swiss demonstrator site manages the flexibility of a pool of buildings, with the goal of contributing to the energy market. All buildings are equipped with heat pumps and are grouped into a single pool to aggregate their flexibility potential. Flexibility is realized via load shedding: FOs are generated with the possibility of temporarily turning off heat pumps and, if the optimization process indicates this action and comfort constraints are respected, the selected heat pumps are switched off. The site includes 63 buildings, and the service has been active for 18 months, from July 2022 to December 2023. The average daily flexibility for the entire pool varies significantly, being less than 40 kWh/day during summer and having its highest values in winter, reaching above 255 kWh/day during January.
The French demonstrator site provides a digital energy coaching service for flexibility. A dedicated application notifies grid users in case of a potential power outage, prompting them to reduce or shift their loads in exchange for a financial incentive. This service has been deployed to 124 individual customers over a 50-day period, arbitrarily selected between March and July 2023, yielding several MWh of flexibility.
The Danish demonstrator site provides services of room-based heating control and cost optimization, by using flexibility provided by heat pumps. This service was deployed for 11 buildings throughout 2023. Results show a 13% cost reduction compared to the baseline, and an average reduction of nearly 20% in energy demand during the three most expensive hours of the day.
For all those demonstrators, dCO contributed making the development of these services easier and better. There are two main reasons for this: i) services can automatically identify the available devices they refer to, as dCO annotates TDs, and are compatible with multiple platforms; ii) dCO supports FOs, which allow for better flexibility representation compared to simple flexibility representations and for better scalability compared to exact flexibility models, as we will see in Section IV-C.
Regarding the integration of this ontology with existing standards or protocols, our ontology allows for better flexibility modeling, being able to capture more complex energy constraints and integrate FOs. Our proposed ontology is also compatible with the Common Information Model (CIM) [^4] https://www.entsoe.eu/digital/common-information-model/, which is the standard from the International Electrotechnical Commission for electric power transmission and distribution, and is capable of representing more complex energy constraints than CIM, which is unable to represent dependency and uncertain energy constraints [CIM]. Finally, SFOs and TECFOs integrate with EEBus [^5] https://www.eebus.org/, another energy data standard that only has simple flexibility models: this is because EEBus builds on SAREF4ENER, which supports slice and total energy constraints.
IV-B Validation
In order to validate dCO, we also used the OntOlogy Pitfall Scanner! (OOPS!) tool [^6] https://oops.linkeddata.es/ to evaluate it. OOPS! [poveda2014oops] is a tool, well-known in the Semantic Web community, designed with the purpose of identifying pitfalls and design errors in ontologies. There are three levels of importance in pitfalls identified by OOPS!. Critical pitfalls need to be resolved, as they may affect the ontology consistency, reasoning, and applicability. Important pitfalls would improve the ontology’s robustness if fixed, but they do not undermine the ontology’s function. Minor pitfalls would improve the ontology’s readability if fixed, but do not pose problems for the ontology. OOPS! identified quite a few minor and important pitfalls, but only 11 critical ones. However, the critical pitfalls were all of the type Defining multiple domains or ranges in properties. In our case, this was not only valid, but also necessary for properties to have multiple domains and ranges to capture the complexity of our flexibility scenario. A simple example of this is the dco:hasEnergyConstraintList property, which can have up to one s4ener:EnergyMin attribute and up to one s4ener:EnergyMax attribute. For this reason, we intentionally designed our ontology with this feature, and therefore none of the critical pitfalls identified by the OOPS! tool undermines the function of the ontology, rather, they are exactly as they should be. Thus, the ontology meets the checked requirements.
IV-C Comparison with SAREF4ENER
In Section II-G, we mentioned that SAREF4ENER is an ontology that allows for a simple description of energy flexibility. FOs, on the other hand, is a model that represents flexibility in mathematical terms; despite having two different purposes, a parallel can be drawn between SAREF4ENER and FO representations, in that they refer to energy profiles and time slices. However, this parallel highlights a crucial weakness of SAREF4ENER in describing energy flexibility. Specifically, SAREF4ENER is unable to model many necessary attributes for FOs: in the main FO message, only State and the attributes defining slice and total energy constraints can be mapped to SAREF4ENER. It is therefore possible to map SFO and TECFO constraints into SAREF4ENER. However, as mentioned in Section III-A, dependency constraints and UFOs cannot be mapped to SAREF4ENER. On the other hand, the dCO flexibility module has been built in order to provide full support for all types of FOs. This means not only that dCO can represent flexibility through more advanced constraints, but also that it allows for communication via a FO message. This way, by using dCO, developers have the possibility to design applications able to exploit the whole range of benefits coming from FO modeling, i.e., i) advanced and more accurate flexibility representation, ii) support for more advanced devices, and iii) really fast optimization and (dis)aggregation of flexibility.
@@ IV-C Comparison with SAREF4ENER
Regarding flexibility representation, SAREF4ENER is able at most to accurately represent flexibility for devices that only need this type of constraint. This is the case with wet devices, such as dishwashers and washing machines, as shown in [DBLP:conf/smartgridcomm/PedersenSN18]. These devices have a load with a fixed power/amount profile, which is only shifted in time. However, for other devices, SFOs/TECFOs are outclassed by DFOs in terms of accuracy in flexibility representation, as we can see in previous literature. The ability to retain flexibility for a model is calculated by the economic metric [Lilliu2021], defined as the amount of profit that can be obtained by optimizing flexibility with the objective of cost reduction. Specifically, for the charge and discharge process of batteries, DFOs retain 6161% of the flexibility, against the 3838% from TECFOs and 1010% from SFOs [Lilliu2021]. UFOs can perform even better, retaining up to 66.466.4% of flexibility, which goes up to 86.886.8% of flexibility when imbalance penalties for violating constraints are high (high penalty case) [DBLP:conf/eenergy/LilliuPSN23]. Regarding EVs, flexibility modeling via DFOs is more accurate than SFOs and TECFOs, as slice constraints and TECs are only a subset of DFO constraints, and therefore the amount of flexibility retained by DFOs is higher: yet, in [DBLP:conf/eenergy/LilliuPSN23], we see that UFOs can retain 9292% of flexibility, against the 77.377.3% retained by DFOs. For the high penalty case, UFOs retain 87.587.5% of flexibility, against the 61.861.8% from DFOs, which still outperform SFOs and TECFOs. Finally, regarding heat pumps, the comparison between DFOs and SFOs has been done in [DBLP:conf/eenergy/SiksnysP16], showing that DFOs can retain more than three times the amount of flexibility from SFOs: in [DBLP:conf/eenergy/LilliuPS23] DFOs for heat pumps have been further improved, showing that they can retain up to 98.998.9% of flexibility according to the economic metric. We show these results in Table III. The ability for dCO to model the significant amounts of flexibility for these devices ensures that developers can create applications that allow for an accurate description of flexibility for batteries, EVs and heat pumps. Table IV summarizes this, showing which use cases are supported by the dCO flexibility module and SAREF4ENER, respectively.
Device SFO TECFO DFO UFO
Battery 10% 38% 61%
66.4%
(86.8%)
EV N/A N/A
77.3%
(61.8%)
92%
(87.5%)
Heat pumps N/A N/A 98.9% N/A
TABLE III: Amount of retained flexibility (in parenthesis, the high penalty case).
@@
TABLE IV: Use cases supported.
Furthermore, FOs are extended over time to represent more use cases and flexibility properties: an example is their capability of modeling different energy vectors from electricity, such as heat [DBLP:conf/eenergy/LilliuPS23]. This has been our rationale to extend SAREF4ENER to represent both FOs in general and, more specifically, DFO and UFO constraints: the resulting ontology can represent all the advanced use cases presented in this paper, as well as provide a solid foundation for future ontology extensions based on new types of FOs.
Therefore, we have shown two important facts. First, DFOs and UFOs vastly outperform SFOs and TECFOs in terms of accuracy of flexibility representation. As it is possible to map only SFOs and TECFOs to SAREF4ENER, the flexibility representation allowed by dCO is much more accurate than the flexibility representation allowed by SAREF4ENER. Second, as shown in Table IV, dCO allows for accurately representing many more types of devices compared to SAREF4ENER: specifically, batteries, EVs and heat pumps can be accurately represented by dCO, but not by SAREF4ENER.
Lastly, dCO allows for support for the fast optimization, aggregation and disaggregation of energy loads provided by FOs. Specifically, FOs allow to aggregate, optimize and disaggregate up to $2\cdot 10^{6}$ of loads for 96 time units time horizons [DBLP:conf/eenergy/LilliuPS23]. In comparison, an exact approach described in [DBLP:conf/eenergy/LilliuPS23] can only optimize up to 500 devices for that time horizon, while other known exact approaches cited in [DBLP:conf/eenergy/LilliuPS23] fail already to perform optimization for more than 6 time units.
@@ V Conclusions and Future Work
-A FlexOffer Message
Here, we describe FO messages and their attributes. Table V shows a full description of an FO message, outlining which attributes are present, whether they are mandatory or not, their data type, and a short description of each attribute. Examples of the message can be seen in Appendix -B onwards: the core part of the message is described first, while the part of the message that describes the constraints changes depending on which constraints are used, and is shown in Appendix -B through -E.
Attribute Mandatory Type Description
ID Yes Integer The ID that identifies the FO
State Yes String
State of the FO
(initial/offered/accepted/rejected/assigned/executed/ invalid/ canceled)
StateReason No String Reason for FO rejection (if State == rejected)
NumSecondsPerInterval Yes Integer Duration (in seconds) of a time interval. The default value is 900
CreationTime Yes Datetime Absolute time at which the FO has been created
CreationInterval No Integer
FO creation interval calculated as epoch value for the
CreationTime / NumSecondsPerInterval ratio
OfferedByID Yes String ID of the FO owner
LocationID No String ID for representing the location of the FO in the grid system
AcceptBeforeTime No Datetime Absolute time before which FO must be accepted
AcceptBeforeInterval No Integer Interval before which FO must be accepted
AssignmentBeforeTime No Datetime Absolute time before which FO must be scheduled
AssignmentBeforeInterval No Integer Interval before which FO must be scheduled
StartAfterTime Yes Datetime Absolute time after which FO must be started
StartAfterInterval No Integer Interval after which FO must be started
StartBeforeTime Yes Datetime Absolute time before which FO must be started
StartBeforeInterval No Integer Interval before which FO must be started
EndAfterTime No Datetime Absolute time after which FO execution must end
EndAfterInterval No Integer Interval after which FO execution must end
EndBeforeTime No Datetime Absolute time before which FO execution must end
EndBeforeInterval No Integer Interval before which FO execution must end
FlexOfferProfileConstraints Yes FlexOfferSlice Constraints for FO profile
FlexOfferPriceConstraint No PriceSlice Constraints for FO price
DefaultSchedule No ScheduleSlice Default energy consumption and time schedule of the FO
FlexOfferSchedule No ScheduleSlice Schedule for the FO
FlexOfferProfileConstraints attributes
EnergyConstraintsList Yes Object
Contains the list(s) of energy constraints for one time unit, expressed
in kWh. Has two sub-elements: lower, and upper.
PriceConstraint No Object
Contains the list(s) of price constraints for one time unit.
Has two sub-elements : lower, and upper.
MinDuration No Integer Minimum duration of the FO
MaxDuration No Integer Maximum duration of the FO
TotalEnergyConstraint No Float List of total energy constraints. Has two sub-elements : lower, and upper.
TotalCostConstraint No Float List of total cost constraints.
ScheduleSlices attributes
Duration No Integer Indicates the duration of the considered slices, in time units.
EnergyAmount Yes Float Indicates the energy consumption for those slices.
Price No Float Indicates the price amount for that slice.
TABLE V: FlexOffer message attributes.
The attributes shown in Table V compose a message: in the following, we show an example for that.
"flexOffer": {
"id": "4188a132-a937-4639-96cf-d8529fa78b86",
"state": "Adaptation",
"stateReason": "FlexOffer Initialized",
"creationInterval": 1726911,
"offeredById": "harry@80060B5E0FD671D58243CE7162A6054719822955",
"locationId": {
"userLocation": {"longitude": 9.990595, "latitude": 57.012293}
"acceptanceBeforeInterval": 1726914,
"assignmentBeforeInterval": 1726914,
"startAfterInterval": 1726912,
"startBeforeInterval": 1726920,
"assignment": "obligatory",
"flexOfferProfileConstraints": $\ldots$\$}
"acceptanceBeforeTime": "2019-04-02T16:30:00.000+0000",
"assignmentBeforeTime": "2019-04-02T16:30:00.000+0000",
"numSecondsPerInterval": 900,
"startAfterTime": "2019-04-02T16:00:00.000+0000",
"startBeforeTime": "2019-04-02T18:00:00.000+0000",
"creationTime": "2019-04-02T15:45:00.000+0000",
"correct": true
Here, we omitted the element FlexOfferProfileConstraints, which describes the constraints defining the FO: its sub-elements are described in the second half of Table V.
-B SFO
Below we can see an example of how the FlexOfferProfileConstraints attribute would appear in the case of the SFO produced for the running example, that we can see in Figure 2(a). In the case we are showing there are also price constraints, where the minimum price for the considered time units is 0.03€, and the maximum is 0.15€.
"flexOfferProfileConstraints": [{
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
Data in ScheduleSlice format has the attributes described in Table V, in the bottom part.
We can also see that Figure 2(a) contains a schedule. This is how it would be encoded in an FO message, assuming this is the default schedule:
"defaultSchedule": {
"scheduleId": 0,
"updateId": 0,
"scheduleSlices": [{"duration": 1, "energyAmount": 0.423, "price": 0.05},
{"duration": 1, "energyAmount": 0.403, "price": 0.1},
{"duration": 1, "energyAmount": 0.388, "price": 0.1},
{"duration": 1, "energyAmount": 0.433, "price": 0.03},
{"duration": 1, "energyAmount": 0.353, "price": 0.03},
{"duration": 1, "energyAmount": 0.393, "price": 0.05},
{"duration": 1, "energyAmount": 0.433, "price": 0.07},
{"duration": 1, "energyAmount": 0.413, "price": 0.07}],
"startTime": "2019-04-02T00:00:00.000+0000"
-C TECFO
For TECFOs the FO message looks the same as for SFOs, except there is an additional TEC at the end of it. Specifically, consider the TECFO shown in Figure 2(b), relative to the running example. The flexOfferProfileConstraints attribute would appear like this:
"flexOfferProfileConstraints": [{
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"energyConstraintList": [{"lower": 0.303, "upper": 0.478}],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
{"TotalEnergyConstraints": [{"lower": 2.592, "upper": 3.381}]}
with the slice constraints listed for each time unit, and the total energy constraint at the end.
-D DFO
The flexOfferProfileConstraints attribute for DFOs looks different from SFOs and TECFOs, as it utilizes the H-representation of the DFO and therefore encodes matrices, instead of lower and upper bounds. An example for the first four time units of the running example, shown in Figure 2(c), can be seen below:
"flexOfferProfileConstraints": [{
"DependencyEnergy ConstraintList": [[0, 1, 0.392],[0, -1, -0.324]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"DependencyEnergy ConstraintList": [[-1, 0, -0.324],[1, 0, 0.392],[-0.221, -1, -0.396],[0.221, 1, 0.514]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"DependencyEnergy ConstraintList": [[0, -1, -0.309],[0, 1, 0.442],[-1, 0, -0.648],[1, 0, 0.819],[-0.127, -1, -0.406],[0.127, 1, 0.531]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"DependencyEnergy ConstraintList": [[0, -1, -0.309],[0, 1, 0.442],[-1, 0, -0.972],[1, 0, 1.246],[-0.088, -1, -0.41],[0.088, 1, 0.537]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
At each time, the DependencyEnergyConstraintList is the representation of a matrix, represented by its rows. Specifically, the one at $t=4$ is the same matrix we described in Section III-C.
-E UFO
Finally, UFO constraints are different from other FO constraints, as they are functions. However, they can be represented by sets of polynomials, as shown in Section III-C: each polynomial is represented by its coefficients, written in increasing order of degree. We can see here the representation of the UFO shown in Figure 5:
"flexOfferProfileConstraints": [{
"UncertainEnergy ConstraintList": [[1]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1
"DependencyEnergy ConstraintList": [[1],[-20.6,66.67],[29.467,-66.67]],
"priceConstraint": {"minPrice": 0.03, "maxPrice": 0.15},
"minDuration": 1,
"maxDuration": 1